### v1

# Generalizing Scaling Laws for Dense and Sparse Large Language Models

Abstract
Over the past few years, the size of language models has grown exponentially, as has the computational cost to train these large models. This rapid growth has motivated researchers to develop new techniques aimed at enhancing the efficiency of the training process. Despite these advancements, optimally predicting the model size or allocating optimal resources remains a challenge. Several efforts have addressed the challenge by proposing different scaling laws, but almost all of them are architecture-specific (dense or sparse). In this work we revisit existing scaling laws and propose a generalized scaling law to provide a unified framework that is applicable to both dense and sparse large language models. We evaluate and compare our proposed scaling law with existing scaling laws to demonstrate its effectiveness.
Index Terms:
LLMs, Dense models, Sparse models, Scaling laws, Sparsity
I Introduction
In recent years, transformer architectures [1] have revolutionized the deep learning approach. These architectures are now the foundation for the majority of popular large language models (LLMs). The reason for the widespread adoption of these architectures is the attention mechanism, which captures the deep contextual relationship in the text data. Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive. As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research. Several approaches have been explored to study the scaling behavior of transformer models. The scaling laws allow researchers and developers to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training. Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse). In this work, we generalize the existing empirical scaling laws and propose a generalized scaling law for both dense and sparse models that enables accurate performance prediction and optimal resource allocation across different architectures.
The field of natural language processing has undergone significant advancement with the development of transformer architectures [1] such as BERT [9] and GPT-3 [10]. Since ChatGPT was publicly released and made available in the fall of 2022, LLMs have gained more attention across diverse domains as a key innovation technology. Several models [11, 12, 13, 14, 15] have been developed over the past few years with traditional dense architecture. A variation of the transformer architecture where a fraction of available parameters are activated—namely, sparse transformers[16]—has become popular lately, and several models [17, 18, 19, 20, 21, 22, 23, 24] have been developed using the concept of sparsification. One common factor among all these models is that as they are evolving: their size have been increasing significantly from GPT-2 with $1.5$ billion parameters [25] to PaLM with 540 billion parameters [14], to DeepSeek-V3, a strong mixture-of-experts (MoE) language model with 671 billion parameters, 37 billion of which are activated for each token [26], which has 94.49% sparsity. With the increase of size, these models require large training datasets, eventually increasing the compute budget [2, 3] to several petaFLOP/s-days of computation to achieve optimal performance, thus making the models expensive to train.
In traditional transformer architectures, parameters of every layer are active for every token. Each token flows through the same feed-forward network; attention layers analyze the entire input sequence simultaneously, and every attention head considers every other position. These kinds of models are dense. The complex and exhaustive architecture allows the dense models to understand the rich contextual meaning of the input dataset. But it comes with high computational cost and scales up with both model size and sequence length.
One way to minimize the cost is to introduce sparsity in a model, where not all the parameters available in the model are used to process a token. Sparsity can be introduced in several ways. We focus on two methods: pruning [27] and mixture of experts [28]. Pruning-based sparsity permanently removes parameters from the model. On the other hand, MoE introduces new experts and uses only a fraction of them to process a token. (A more detailed explanation of pruning and MoE is given in Subsection II-B.)
Foundation models have gained attention across diverse domains because of their high accuracy, reasoning abilities, and generative power. Achieving these capabilities, however, requires vast amounts of training data and substantial compute resources. As a result, optimal allocation of compute, model size, and data volume has become critical to minimizing training costs.
Scaling laws—empirical relationships often expressed as power laws—describe a model’s performance based on allocated resources. These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance. Several such empirical scaling laws based on different LLM architectures have been proposed in recent years.[2, 3, 4, 5, 29, 30, 7, 8, 31] Others are based on training precision [30], distillation [32], and quantization [29]. Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35]. This interest in scaling laws for different domains makes them pivotal in deep neural network research.
In this paper we focus on the scaling laws for dense and sparse LLMs. We make the following contributions:
  • We discuss the existing empirical scaling laws for dense and sparse LLMs and their limitations (Section II).
  • We propose a generalized scaling law for both dense and sparse LLMs and discusses the relationship between the proposed scaling law and the existing laws (Section III).
  • We evaluate and compare the performance of the proposed scaling law and the existing laws using some test models to demonstrate its effectiveness, and we present hyperparameter optimization results for some existing scaling laws (Section IV).
Section V summarizes this work and suggests avenues for potential future research.
II Scaling Laws for LLMs
Transformer models use a fully connected attention mechanism in which the token is addressed by the other tokens of the sequence. This process, which activates all the parameters in the model, is computationally expensive. The cost of training a large transformer model depends on several factors. Based on the number of parameters available in the architecture and the number of parameters that are being used during pretraining, transformer models can be categorized into dense and sparse transformers, which have shown different scaling behaviors. In the following, we discuss the scaling laws for dense and sparse LLMs.
II-A Dense Scaling Laws
Traditional transformer models are dense in nature. Kaplan et al. [2] studied the behavior of LLMs with a series of experiments and different hyperparameters: model size, model shape, total compute, data volume, and batch size. The authors observed that performance or model loss during pretraining follows a power law relationship with the total number of parameters, the total number of training tokens, and the total compute budget. They established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training. It also depends mildly on other architectural hyperparameters such as model shape: depth and width. For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [2] because approximately 6 floating point operations are needed for every parameter in the model and every training sample. They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times. They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
$L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}.$ (1)
The best-fitted values for the coefficients and exponents in Equation 1 are listed in Table I.
Coefficients $\alpha_{N}$ $\alpha_{D}$ $N_{C}$ $D_{C}$
Values 0.076 0.103 $6.4\times 10^{13}$ $1.8\times 10^{13}$
TABLE I: Fitted parameter values for Equation 1
Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.  [3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion. In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well. They represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps. Based on their observations, the authors proposed the following scaling law:
$L(N,D)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (2)
This law establishes a power-law relationship between loss $L$, number of parameters $N$, and the data size $D$. Here $e$ captures the entropy of the natural text, which is the minimum loss as both $N$ and $D$ approach infinity. The second term (parameter-dependent term) captures the inverse relationship between model size $N$ and loss. The third term (data-dependent term) captures the impact of the data size $D$ on loss. The constants $a$ and $b$ and the exponents $\alpha$ and $\beta$ are determined empirically through some experiments and fitting the data. Its best-fitted values for the coefficients and exponents are listed in Table II. Equation 2 currently is a well-known empirical scaling law for dense LLMs.
Coefficients $e$ $a$ $b$ $\alpha$ $\beta$
Values $1.69$ $406.4$ $410.7$ $0.34$ $0.28$
TABLE II: Fitted parameter values for Equation 2
II-B Sparse Scaling Laws
Sparse models use only a subset of the parameters available in the models. As a result, Equation 2 needs to be adjusted and extended. Several works had proposed scaling laws for sparse transformer models [4, 5, 8]. These scaling laws are based on different sparsification methods. In this work we focus on two types of sparsifications: pruning and mixture-of-experts (MoE).
II-B1 Pruned Transformer Models
One way to introduce sparsity in transformer models is pruning, where a number of parameters are removed from the architecture and are not used during the forward pass or backpropagation, hence making no contribution to the training process. Pruning can be categorized into three main groups [27]. In structured pruning, entire higher-level units or even whole layers are removed; this process delivers universal speedup but sometimes at the cost of accuracy. In unstructured pruning, individual weights are removed from anywhere in the network; usually the least important weights are pruned, with the importance based on certain criteria. In semi-structured pruning, one tries to strike a balance between unstructured and structured pruning approaches by enforcing fixed pruning patterns in small regions of the network.
Frantar et al. [4] studied the scaling behavior of models pruned by unstructured pruning. Sparsity $S$ is defined as $S=\frac{\text{Total Parameters - Nonzero Parameters}}{\text{Total Parameters}}$. The experimental results with a total of $48$ models trained showed that sparsity introduced by pruning affects only the parameter-dependent term of Equation 2 and has no impact on the data-dependent term. The loss was fitted as a function of nonzero parameters $N$, data $D$, and sparsity $S$ as shown in the following equation.
$L(N,D,S)=\left(a_{S}(1-S)^{b_{S}}+c_{S}\right).\left(\frac{1}{N}\right)^{b_{N}}+\left(\frac{a_{D}}{D}\right)^{b_{D}}+c$ (3)
Following the similar approach of [3] to find the optimal values for the coefficients of Equation 3, the best-fitted values for these coefficients are determined empirically through some experiments and fitting the data shown in Table III.
Coefficients $a_{S}$ $b_{S}$ $c_{S}$ $b_{N}$ $a_{D}$ $b_{D}$ $c$
Values $16.8$ $0.722$ $45$ $0.245$ $6.90\times 10^{8}$ $0.203$ $0.651$
TABLE III: Fitted parameter values for Equation 3
To draw the similarities between Equations 2 and 3, we used the similar symbols in Equation 2 to reformat the coefficients in Equation 3. We used $e$ instead of $c$ to capture the lower bound on the loss that represents the inherent stochasticity of the modeling problem. We replaced $b_{N}$ and $b_{D}$ with $\alpha$ and $\beta$, respectively, and ${a_{D}}^{b_{D}}$ with $b$. The result is the following equation:
$L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (4)
Equation 4 helps visualize the transition from the dense scaling law Equation 2 to the sparse scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity. Based on the changes in the coefficient representation, the best-fitted values need to be updated correspondingly. The updated values for the coefficients in Equation 4 are calculated and listed in Table IV.
Coefficients $a_{S}$ $b_{S}$ $c_{S}$ $b$ $\alpha$ $\beta$ $e$
Values $16.8$ $0.722$ $45$ $62.271$ $0.245$ $0.203$ $0.651$
TABLE IV: Fitted parameter values for Equation 4
While Equation 3 functionally captures the effect of sparsity on loss, the maximum sparsity used in the experiments was $87.5\%$ [4]. We have generated an estimated IsoFLOP dataset for the compute budget $1e20$ from the experiments conducted in [5], which included sparsities of more than $87.5\%$. When we used the new dataset to plot the scaling behavior of Equation 3, we noticed a spike for smaller nonzero parameters, as shown in Figure 1, where each line connects models with different sparsity and model sizes trained under the same amount of compute budget $1e20$ with nonzero parameters as x-axis and training loss as y-axis. Increasing the nonzero parameters results in decreasing the data size. However, based on the work in [4, 5], sparsity should provide a smooth performance gain as we increase sparsity. This indicates that Equation 3 does not capture the scaling behavior of highly sparse models very well.
Figure 1: IsoFLOP plot of Equation 3 for different sparsities
On the other hand, the largest model used in [4] had only $85M$ nonzero parameters. Compared with the current work, these models are relatively much smaller. This work used the scaling law Equation 2 [3] as the base for the approach. For $0\%$ sparsity, or for dense models, the loss should be equal to Equation 2. But for the same data set with no sparsity, Equation 3 provides different coefficient values from those of Equation 2, as shown in Figure 2, there is a difference of up to 2.5 in loss. Thus, we can conclude that Equation 2 is not a special case of Equation 3 if the sparsity is 0.
Figure 2: Comparison between Equations 2 and 3 for dense models
II-B2 MoE Models
Another way to introduce sparsity is to use the mixture-of-experts method [28] where the feed-forward network for a transformer model is replicated multiple times and each of these copies is called an “expert.” Only a fraction of experts are activated per token, and this number is a fixed value. The tokens are routed to different experts during the pretraining process. The size of the experts is usually equal to the size of the feed-forward network.
Abnar et al. [5] investigated the sparse MoEs through a large-scale empirical study. The sparsity $S$ was defined as $S=\frac{E-K}{E}$, where $E$ is the total number of experts and $K$ is the number of active experts. For five different compute budgets and seven different sparsities, different models were trained. Based on the experimental results, the authors proposed the following scaling law:
$L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}.$ (5)
Through a grid initialization, for a sparsity of $98\%$, the best-fitted coefficient values in Equation 5 are determined empirically through the experiments and fitting the data shown in Table V.
Coefficients $e$ $a$ $b$ $c$ $d$
Values $0.94$ $16612.50$ $5455.67$ $0.4598$ $17.26$
Coefficients $\alpha$ $\beta$ $\lambda$ $\delta$ $\gamma$
Values $0.5962$ $0.3954$ 0.1666-0.1666 $0.1603$ $0.1595$
TABLE V: Fitted parameter values for Equation 4
Equation 5 introduced two different terms in the scaling law along with the previous three terms from [3] to capture the impact of sparsity. The newly introduced terms are interdependent. The fifth term in Equation 5 can be interpreted as an active parameter term because the denominator term is a multiplication of $N$ and 1S1-S. Because the total number of parameters already includes the active parameters, however, having two different terms for total parameters and active parameters in Equation 5 causes some redundancy in the loss function.
This work also used Equation 2 as the base. But for $0\%$ sparsity we apply the models from [5] to Equations 5 and 2 to generate Figure 3 and notice a difference in loss values. This indicates that Equation 2 is not a special case of Equation 5 if the sparsity is 0.
Figure 3: Comparison between Equations 2 and 5 for dense models
III Generalizing Scaling Laws
We observed that the previous work and their scaling laws are based on the number of nonzero or active parameters. Dense models use all the parameters, and hence in [3] $N$ was used as the total number of parameters in its scaling law. In [4] the nonzero parameters $N$ were considered in its scaling law 3. In [5] the authors modified the compute budget $C=6ND$ from [2] to $C=6N_{a}D$, where $N_{a}$ is the number of active parameters. These scaling laws do not cover the dense scaling law when the sparsity is 0.
These observations and limitations motivate us to generalize the scaling laws of [3, 4, 5] into a single representation that can effectively capture the experimental behavior and features observed in these works. We want to make sure that the new scaling law works for both dense and sparse models. As we increase model size, data size, and compute, we want to get better performance [2, 3, 4, 5]. Similarly, as we increase the sparsity, we want to obtain performance improvement [4, 5]. Furthermore, for $0\%$ sparsity or dense models, we need to make sure that our scaling law behaves the same as the original scaling law in Equation 2.
For simplicity, we use the term active parameters $N$ for active/nonzero parameters in the MoE and pruned models. For the dense models, the total number of active parameters is the total number of parameters. We use Equation 2 as the base to extend it to a generalized scaling law. We define the sparsity as follows:
$S=\frac{\text{Total Parameters - Active Parameters}}{\text{Total Parameters}}$ (6)
Where $0\leq S<1$ because the active parameters in a model cannot be zero. Therefore $\frac{N}{1-S}$ is the total number of parameters. In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$. Similar to the approach in [4] we introduce an upper-bound constant sparsity factor $c$, which is caused by the sparsity; but this factor, which contributes to the loss, depends on the sparsity $S$. So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$. For sparsity $S=0$ or dense models, this term is the same as the original parameter-dependent term in Equation 2.
We have noticed that different approaches adjusted the entropy $e$ in different ways, as shown in Tables II, IV, and V. The value of $e$ is impacted by the sparsity. We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$. Accumulating these changes, we propose the following scaling law:
$L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (7)
This equation quantifies the impact of sparsity on loss. Notice that we did not change the values of the data-dependent term in Equation 2, because the work in [4, 5] has shown experimentally that sparsity has no effect on this term.
The fitted coefficient values in our proposed Equation 7 are listed in Table VI. Since we have used the empirical scaling law in [3] as the base, we simply apply the coefficient values in Table II to Equation 7. For the sparsity factor $c$, we use the dataset for the largest model with the largest sparsity in [4] to fit Equation 7 to obtain the value for the sparsity factor $c$. The value of $\gamma$ was empirically obtained.
Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
Values $1.69$ $406.4$ $410.7$ $93.45$ $0.34$ $0.28$ $10^{-2}$
TABLE VI: Fitted parameter values for law 7
Notice that Equation 7 is similar to Equation 4; however, the coefficients and the parameter-dependent term are different. If the sparsity is 0, Equation 4 is not the same as Equation 2, as shown in Figure 2.
Lemma: If the sparsity $S$ is 0, Equation 7 is the same as Equation 2.
If the sparsity $S$ is 0, the total number of parameters is equal to the total number of active parameters. From Equation 7 we have
$L(N,D,0)=e(1-0)^{\gamma}+\left(a(1-0)^{\alpha}+c\cdot 0\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (8)
Then, simplifying Equation 8, we have
$L(N,D,0)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}.$
From Equation 2, we have
$L(N,D)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}.$
Because the coefficients $e,a,b,\alpha,\beta$ are the same, we have
L(N,D,0)L(N,D)=0L(N,D,0)-L(N,D)=0.
IV Performance Results and Discussion
IV-A Performance Results
In this section we compare the performance between our proposed scaling law and the previous scaling laws for dense and sparse models. We have collected the experimental datasets (number of parameters, number of tokens, and sparsity) used in [3, 4, 5]. Hoffman et al. [3] provided a list of nine models with number of parameters and training tokens used in their experiments. Franta et al.  [4] provided four different numbers of nonzero parameters, three different numbers of training tokens, and four different sparsity levels with a total of $48$ experimental runs to collect their experimental results. Abnar et al.  [5] used five different numbers of compute budgets, seven different sparsity levels, and a range of model sizes. Using the formula $C=6ND$, we calculated their number of training tokens and listed the ranges in Table VII. We use these experimental datasets to generate plots for their respective scaling laws, and we use the same datasets to generate plots using our proposed scaling law in order to do a head-to-head comparison.
Scaling Law Parameters Tokens Sparsities(%) Test Models
Hoffman et al. [3] 400M10T400M{-}10T 8B216.2B8B{-}216.2B - 99
Frantar et al. [4] 1.3M85M1.3M{-}85M 16B65B16B{-}65B $0,50,75,87.5$ $48$
Abnar et al. [5] 329M21.2B329M{-}21.2B 15B128B15B{-}128B $0,25,50,75,90,95,98$ $35$
TABLE VII: Parameters, tokens, sparsity, and number of test models collected from the corresponding works.
When we compare our proposed scaling law in Equation 7 with the dense scaling law [3] in Equation 2, as we discussed in the lemma in the preceding section, if the sparsity is 0, Equation 7 is the same as Equation 2. Figure 4 shows the loss vs compute budget using the dataset in [3]. On the $x$-axis is compute. which depends on parameters $N$ and tokens $D$ and hence captures the impact of both ($C=6ND$). On the $y$-axis is the calculated loss. We see an identical plot, indicating that our proposed scaling law behaves the same as the original scaling law for dense models.
Figure 4: Comparison between Hoffman’s scaling law in Equation 2 and the proposed scaling law in Equation 7
We compare the proposed scaling law in Equation 7 with the scaling law [4] in Equation 3. We use the 48 models to conduct the experiments in [4] for performance comparison. Figures 5 (a) and 5 (b) show that our proposed scaling law replicates the behavior of the scaling law in [4]. We observe that with the increase of the sparsity the loss decreases. However, there is a difference between the loss values for both scaling laws which is similar to Figure 2. This is because the models used for the experiments in [4] are relatively much smaller (the maximum number of parameters just 85M) than other works, and the adjusted coefficient values in the proposed scaling law are for larger models, as shown in Figure 4 where the loss decreases with increasing the model size and compute budget. Hence we see a relatively larger loss for the much smaller datasets.
(a) Scaling behavior for Equation 3 with varying compute budget
(b) Scaling behavior for Equation 7 with varying compute budget
Figure 5: Comparison between (a) the scaling law in Equation 3 and (b) the proposed scaling law in Equation 7
We also compare our proposed scaling law in Equation 7 with the scaling law [5] in Equation 5. For performance comparison we use the interpolated values of parameters, tokens, and the exact compute budget of the 35 models the authors used to conduct the experiments in [5]. We have drawn compute budget vs loss plots in Figures 6(a) and 6(b) to compare Equations 5 and 7. We observe that with the increase of the sparsity and compute, the loss decreases; and both figures show a similar pattern. Increasing the compute budget means an increase in the active parameter size ($N_{a}$) and number of tokens ($D$). This indicates that the proposed scaling law shows scaling behavior similar to that in [5].
(a) Scaling behavior for Equation 5 with varying compute budget
(b) Scaling behavior for Equation 7 with varying compute budget
Figure 6: Comparison between (a) the scaling law in Equation 5 and (b) the proposed scaling law in Equation 7
IV-B Hyperparameter Optimization for Better Coefficients
Most of the previous scaling laws discussed in this paper used grid search to find the best values for the coefficients. While grid search is a commonly used approach, it is still inefficient. As an alternative, we have used ytopt [36, 37, 38] for hyperparameter optimization. ytopt is a Bayesian optimization–based autotuner that builds a surrogate model to explore the promising regions to find the optimal values with fewer evaluations. In our previous work [39], ytopt outperformed grid search, random search, a genetic algorithm, and XGBoost in accuracy and tuning time. Unlike traditional autotuners, ytopt leverages libEnsemble’s [40] asynchronous manager-worker framework to run multiple evaluations in parallel, reducing the one-by-one autotuning time.
To use ytopt for the optimization process for finding the best coefficients for the previous scaling laws, we use the best coefficients for the scaling laws in Tables II, IV, and V as the baseline and the default values to define the proper search space for each coefficient. We carefully define the range for each coefficient for Equations 2, 3, and 5 and then use ytopt to run the optimization process. The generated hyperparameter optimization results in Figure 7(a), (b), and (c) illustrate that ytopt identifies the best coefficient configuration (combination of coefficients) over time (blue) with the baseline (red) for the default values and that using ytopt results in lower loss values than the coefficient-based loss values (red, baseline) using grid search.
(a) Tuning Loss for Equation 2
(b) Tuning Loss for Equation 3
(c) Tuning Loss for Equation 5
Figure 7: Loss calculated by the coefficients generated by ytopt and grid search. The red lines are the loss using the best coefficients provided in the respective work, and the blue dots are the loss tuned for different evaluations over time using ytopt.
IV-C Discussion
In summary, we observe that our proposed scaling law Equation 7 effectively captures the scaling behavior of Equations 2, 3, and 5 for dense and sparse LLMs. In Figure 8 the proposed scaling law shows no spike for higher sparsity with the fixed compute budget 1e20 as we observed for Equation 3 in Figure 1. This indicates that the proposed scaling law captures the scaling behaviors for higher sparsity. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with the given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt [36, 37].
In Figure 5 we observe that the range of loss values is different for Equations 3 and 7. This difference is mainly because the experiments done for Equation 3 use much smaller model sizes with at most 85M. For the scaling law in Equation 5 we can see that the results are almost identical. These results further confirm that our proposed scaling law captures the scaling behaviors of the MoE models.
Figure 8: IsoFLOP plot of the proposed scaling law in Equation 7 for different sparsities
In this work we have generalized the scaling behaviors of dense and sparse models. Our generalized scaling law extends and covers several existing scaling laws. Its limitations are related to the sparsity factor $c$ and the coefficient $\gamma$, which need to be further optimized for larger models. There are other ways [41, 42] to introduce sparsification in a model. Different training approaches [30, 29] may yield different scaling behavior as well. Further work can be done to incorporate these scaling laws into our generalized scaling law.
V Conclusions
Current large language models result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks. However, these models are costly to train. As LLMs get larger and more expensive, it is crucial for researchers and developers to estimate the expected performance beforehand. Scaling laws thus are critical. But while architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical. Our generalized scaling law covers densely activated, pruned, and MoE models. Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with a given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt.
Because LLMs inference can be memory-bound and compute-intensive, future work will extend the generalized scaling law in training to not only incorporate other scaling laws [30, 32, 29] but also take into account inference [43], especially recent agentic AI workloads for answering complex questions through reasoning using chain-of-thought prompting. Since the behavior of scaling laws varies based on training and inference approaches, these approaches can be incorporated in the generalized scaling law as well. Unifying different scaling laws into one single representation lays the foundation of architecture-agnostic representation of the scaling behavior of LLMs. As LLM models continue to diversify, this generalized approach will allow us to anticipate the return based on their allocated compute resources.
Acknowledgments
This work was supported DOE ASCR RAPIDS2 and OASIS. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.

### v2

@@
 These architectures are now the foundation for the majority of popular large language models (LLMs).
-The reason for the widespread adoption of these architectures is the attention mechanism, which captures the deep contextual relationship in the text data.
+As significantly increasing the model size, dataset size and the amount of compute used for training, LLM performance improves dramatically.
 Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive.
 As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research.
-Several approaches have been explored to study the scaling behavior of transformer models.
-The scaling laws allow researchers and developers to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training.
+Several approaches have been explored to study the scaling behavior of transformer models to propose scaling laws to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training.
 Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse).
@@
 These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance.
-Several such empirical scaling laws based on different LLM architectures have been proposed in recent years.[2, 3, 4, 5, 29, 30, 7, 8, 31] Others are based on training precision [30], distillation [32], and quantization [29].
+Several such empirical scaling laws based on different LLM architectures have been proposed in recent years [2, 3, 4, 5, 6, 7, 8, 29].
+Others are based on training precision [30], distillation [31], and quantization [32].
 Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35].
@@
 For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [2] because approximately 6 floating point operations are needed for every parameter in the model and every training sample.
-They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times.
+They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times to avoid overfitting.
 They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
-$L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}.$ (1)
+$L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}$ (1)
 The best-fitted values for the coefficients and exponents in Equation 1 are listed in Table I.
@@
 Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.
-[3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion.
+[3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion to avoid overfitting.
 In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well.
@@
 The result is the following equation:
-$L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (4)
+$L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (4)
 Equation 4 helps visualize the transition from the dense scaling law Equation 2 to the sparse scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity.
@@
 Based on the experimental results, the authors proposed the following scaling law:
-$L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}.$ (5)
+$L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}$ (5)
 Through a grid initialization, for a sparsity of $98\%$, the best-fitted coefficient values in Equation 5 are determined empirically through the experiments and fitting the data shown in Table V.
@@
 Where $0\leq S<1$ because the active parameters in a model cannot be zero.
-Therefore $\frac{N}{1-S}$ is the total number of parameters.
+When $S=0$, it means that the total parameters is the same as the active parameters.
+Therefore, $\frac{N}{1-S}$ is the total number of parameters.
 In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$.
@@
 The value of $e$ is impacted by the sparsity.
-We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$.
+We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$, where $\gamma$ is a constant.
 Accumulating these changes, we propose the following scaling law:
-$L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (7)
+$L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (7)
 This equation quantifies the impact of sparsity on loss.
@@
 Values $1.69$ $406.4$ $410.7$ $93.45$ $0.34$ $0.28$ $10^{-2}$
-TABLE VI: Fitted parameter values for law 7
+TABLE VI: Fitted parameter values for Equation 7
 Notice that Equation 7 is similar to Equation 4; however, the coefficients and the parameter-dependent term are different.
@@
 If the sparsity $S$ is 0, the total number of parameters is equal to the total number of active parameters. From Equation 7 we have
-$L(N,D,0)=e(1-0)^{\gamma}+\left(a(1-0)^{\alpha}+c\cdot 0\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (8)
+$L(N,D,0)=e(1-0)^{\gamma}+\left(a(1-0)^{\alpha}+c\cdot 0\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (8)
 Then, simplifying Equation 8, we have
@@
 V Conclusions
-Current large language models result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks.
+Current LLMs result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks.
 However, these models are costly to train.

### v3

@@
 Abstract
-Over the past few years, the size of language models has grown exponentially, as has the computational cost to train these large models.
-This rapid growth has motivated researchers to develop new techniques aimed at enhancing the efficiency of the training process.
-Despite these advancements, optimally predicting the model size or allocating optimal resources remains a challenge.
-Several efforts have addressed the challenge by proposing different scaling laws, but almost all of them are architecture-specific (dense or sparse).
-In this work we revisit existing scaling laws and propose a generalized scaling law to provide a unified framework that is applicable to both dense and sparse large language models.
-We evaluate and compare our proposed scaling law with existing scaling laws to demonstrate its effectiveness.
-Index Terms:
-LLMs, Dense models, Sparse models, Scaling laws, Sparsity
+Despite recent advancements of large language models (LLMs), optimally predicting the model size for LLM pretraining or allocating optimal resources still remains a challenge.
+Several efforts have addressed the challenge by proposing different empirical scaling laws, but almost all of them are architecture-specific (dense or sparse).
+In this work we revisit existing empirical scaling laws and propose a generalized scaling law to provide a unified framework that is applicable to both dense and sparse large language models.
+We evaluate and compare our proposed scaling law with existing scaling laws and demonstrate that our proposed scaling law captures the scaling behavior of existing scaling laws.
+Further, we show an IsoFLOP comparison between our proposed scaling law and the state-of-the-art scaling law to illustrate the effectiveness of our proposed scaling law for Mixture-of-Expert (MoE)-based very large LLMs like DeepSeek-V3.
+Our proposed scaling law can be used to estimate the best model hyperparameters (Model size, Tokens and Compute) for a given sparsity or to identify the optimal sparsity for the given model hyperparameters.
 I Introduction
-In recent years, transformer architectures [1] have revolutionized the deep learning approach.
-These architectures are now the foundation for the majority of popular large language models (LLMs).
-As significantly increasing the model size, dataset size and the amount of compute used for training, LLM performance improves dramatically.
-Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive.
-As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research.
-Several approaches have been explored to study the scaling behavior of transformer models to propose scaling laws to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training.
-Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse).
-In this work, we generalize the existing empirical scaling laws and propose a generalized scaling law for both dense and sparse models that enables accurate performance prediction and optimal resource allocation across different architectures.
-The field of natural language processing has undergone significant advancement with the development of transformer architectures [1] such as BERT [9] and GPT-3 [10].
-Since ChatGPT was publicly released and made available in the fall of 2022, LLMs have gained more attention across diverse domains as a key innovation technology.
+In recent years, transformer architectures [37] have revolutionized the deep learning approach especially the field of natural language processing with the development of transformer-based architectures [37] such as BERT [12] and GPT-3 [3] etc.
+Transformer architecture is now the foundation for the majority of popular large language models (LLMs).
+Since the release of ChatGPT LLMs have gained more attention across diverse domains as a key innovative technology.
+With the growing popularity across diverse domain there is an increasing need to scale LLM knowledge and accuracy while also reducing the inference latency and training cost.
+However, achieving these goals is fundamentally challenging due to the computational structure of standard transformer models, as computation grows rapidly with the increase of model size and context length.
+In traditional transformer architecture, parameters in every layer are activated to process every token.
+Each token passes through the same feed-forward network, and in a multi-head attention setting, every attention considers every other position.
+These traditional forms of transformers are called dense transformers.
+The complex and exhaustive architecture allows the dense models to understand the rich contextual meaning of the input dataset, but it comes with a high computational cost that scales with both model size and sequence length.
+One way to reduce the cost is to introduce sparsity where all the available parameters are not used to process a token in a sparse model.
+There are several ways to introduce sparsity, but in this work we focus on two methods: pruning [6] and mixture of experts (MoE)  [33].
+Pruning-based sparsity permanently removes parameters from the model.
+Pruning can be categorized into three main groups [6] namely: structured pruning, unstructured pruning, and semi-structured pruning.
+In structured pruning, entire higher-level units or even whole layers are removed; this process delivers universal speedup but sometimes at the cost of accuracy.
+In unstructured pruning, individual weights are removed from anywhere in the network; usually the least important weights are pruned.
+In semi-structured pruning, one tries to strike a balance between unstructured and structured pruning approaches by enforcing fixed pruning patterns in small regions of the network.
+On the other hand, in MoE-based models the feed-forward network for a transformer model is replicated multiple times, and each of these copies is called an “expert”.
+In the MoE-based models only a fraction of experts are activated per token, instead of all the experts and that’s how sparsity is introduced.
+The number of activated experts is a fixed value and determines the sparsity level.
+The tokens are routed to different experts during the pretraining process.
+The size of the experts is usually equal to the size of the feed-forward network but can differ too.
 Several models [31, 34, 36, 8, 28] have been developed over the past few years with traditional dense architecture.
-A variation of the transformer architecture where a fraction of available parameters are activated—namely, sparse transformers[16]—has become popular lately, and several models [17, 18, 19, 20, 21, 22, 23, 24] have been developed using the concept of sparsification.
-One common factor among all these models is that as they are evolving: their size have been increasing significantly from GPT-2 with $1.5$ billion parameters [25] to PaLM with 540 billion parameters [14], to DeepSeek-V3, a strong mixture-of-experts (MoE) language model with 671 billion parameters, 37 billion of which are activated for each token [26], which has 94.49% sparsity.
-With the increase of size, these models require large training datasets, eventually increasing the compute budget [2, 3] to several petaFLOP/s-days of computation to achieve optimal performance, thus making the models expensive to train.
-In traditional transformer architectures, parameters of every layer are active for every token.
-Each token flows through the same feed-forward network; attention layers analyze the entire input sequence simultaneously, and every attention head considers every other position.
-These kinds of models are dense.
-The complex and exhaustive architecture allows the dense models to understand the rich contextual meaning of the input dataset.
-But it comes with high computational cost and scales up with both model size and sequence length.
-One way to minimize the cost is to introduce sparsity in a model, where not all the parameters available in the model are used to process a token.
-Sparsity can be introduced in several ways.
-We focus on two methods: pruning [27] and mixture of experts [28].
-Pruning-based sparsity permanently removes parameters from the model.
-On the other hand, MoE introduces new experts and uses only a fraction of them to process a token.
-(A more detailed explanation of pruning and MoE is given in Subsection II-B.)
-Foundation models have gained attention across diverse domains because of their high accuracy, reasoning abilities, and generative power.
-Achieving these capabilities, however, requires vast amounts of training data and substantial compute resources.
-As a result, optimal allocation of compute, model size, and data volume has become critical to minimizing training costs.
-Scaling laws—empirical relationships often expressed as power laws—describe a model’s performance based on allocated resources.
+But lately, sparse transformers[7]—have become extremely popular, and several models [14, 15, 43, 27, 25, 10, 22, 18] have been developed using the concept of sparsification.
+One common factor among all these models is that as they are evolving, their size has been increasing significantly from GPT-2 with $1.5$ billion parameters [30] to PaLM with 540 billion parameters [8], to DeepSeek-V3, a strong mixture-of-experts (MoE) language model with 671 billion parameters, 37 billion of which are activated for each token [11], which has 94.49% sparsity.
+The reason is, it has been proven that as model size, dataset size, and the amount of compute used for training increase, LLM performance improves smoothly [23].
+However with the increase of size, these models require large training datasets, eventually increasing the compute budget [23, 20] to several petaFLOP/s-days of computation to achieve optimal performance.
+This makes LLM training time-consuming and resource intensive.
+As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training is crucial.
+Scaling laws are the empirical relationships, often expressed as power laws—describe a model’s performance based on allocated resources.
 These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance.
-Several such empirical scaling laws based on different LLM architectures have been proposed in recent years [2, 3, 4, 5, 6, 7, 8, 29].
-Others are based on training precision [30], distillation [31], and quantization [32].
-Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35].
+Several empirical scaling laws based on different LLM architectures have been proposed in recent years [23, 20, 16, 1, 19, 9, 26, 17].
+Few scaling law works focused on training precision [24], distillation [4], and quantization [5].
+Several studies have also been done on vision [16], audio [13, 29], and mixed-modal language models [2].
 This interest in scaling laws for different domains makes them pivotal in deep neural network research.
-In this paper we focus on the scaling laws for dense and sparse LLMs. We make the following contributions:
+However, until now, nearly all existing approaches have been architecture-specific (dense or sparse) and also use different scaling law coefficients (details in Section II) which makes it difficult to perform cross-architecture comparisons and budget planning.
+In this work we show that transformers, regardless of dense and sparse architectures, exhibit a common scaling behavior which depends on the number of active/nonzero parameters.
+To demonstrate that, we generalize the existing empirical scaling laws to propose a scaling law that enables accurate performance predictions and optimal resource allocation across architectures.
+In this paper we focus on the scaling laws for both dense and sparse LLMs.
+We make the following contributions:
   • We discuss the existing empirical scaling laws for dense and sparse LLMs and their limitations (Section II).
-  • We propose a generalized scaling law for both dense and sparse LLMs and discusses the relationship between the proposed scaling law and the existing laws (Section III).
+  • We propose a generalized scaling law for both dense and sparse LLMs and discuss the relationship between the proposed scaling law and the existing laws (Section III).
   • We evaluate and compare the performance of the proposed scaling law and the existing laws using some test models to demonstrate its effectiveness, and we present hyperparameter optimization results for some existing scaling laws (Section IV).
@@
 II Scaling Laws for LLMs
-Transformer models use a fully connected attention mechanism in which the token is addressed by the other tokens of the sequence.
-This process, which activates all the parameters in the model, is computationally expensive.
-The cost of training a large transformer model depends on several factors.
-Based on the number of parameters available in the architecture and the number of parameters that are being used during pretraining, transformer models can be categorized into dense and sparse transformers, which have shown different scaling behaviors.
-In the following, we discuss the scaling laws for dense and sparse LLMs.
-II-A Dense Scaling Laws
+Based on the number of parameters available in the architecture and the number of parameters that are being used during pretraining, transformer models can be categorized into dense and sparse transformers.
+Previous works [31, 34, 36, 8, 28, 14, 15, 43, 27, 25, 10, 22, 18] have shown that these models exhibit different behavior during pretraining as they scale, and hence their scaling behaviors are studied separately.
+II-A Scaling Laws for Dense Models
 Traditional transformer models are dense in nature.
 Kaplan et al.
-[2] studied the behavior of LLMs with a series of experiments and different hyperparameters: model size, model shape, total compute, data volume, and batch size.
-The authors observed that performance or model loss during pretraining follows a power law relationship with the total number of parameters, the total number of training tokens, and the total compute budget.
-They established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training.
+[23] studied the behavior of dense LLMs with a series of experiments by monitoring different hyperparameters: model size, model shape, total compute, data volume, and batch size during pretraining.
+The authors established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training.
 It also depends mildly on other architectural hyperparameters such as model shape: depth and width.
 For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [23] because approximately 6 floating point operations are needed for every parameter in the model and every training sample.
-They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times to avoid overfitting.
-They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
+The authors also suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times to avoid overfitting.
+They represented training loss as a function of number of parameters and data size and proposed the empirical scaling law:
 $L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}$ (1)
@@
 Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.
-[3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion to avoid overfitting.
-In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well.
-They represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps.
-Based on their observations, the authors proposed the following scaling law:
+[20] extended and corrected Equation 1 in [23] and concluded that model size and number of tokens should be increased in proportion to avoid overfitting.
+The authors represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps.
+Based on the observations from their three experiments, the authors proposed Equation 2 as their scaling law.
 $L(N,D)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (2)
-This law establishes a power-law relationship between loss $L$, number of parameters $N$, and the data size $D$.
+Figure 1: Comparison between Hoffman scaling law and Frantar scaling law at $0\%$ sparsity.
+This law establishes a power-law relationship between loss $L$, NN, and $D$.
 Here $e$ captures the entropy of the natural text, which is the minimum loss as both $N$ and $D$ approach infinity.
-The second term (parameter-dependent term) captures the inverse relationship between model size $N$ and loss.
-The third term (data-dependent term) captures the impact of the data size $D$ on loss.
+The second term (parameter-dependent term) captures the inverse relationship between $N$ and loss.
+The third term (data-dependent term) captures the impact of $D$ on loss.
 The constants $a$ and $b$ and the exponents $\alpha$ and $\beta$ are determined empirically through some experiments and fitting the data.
 Its best-fitted values for the coefficients and exponents are listed in Table II.
-Equation 2 currently is a well-known empirical scaling law for dense LLMs.
+Equation 2 currently is the state-of-the-art empirical scaling law for dense LLMs.
+For the rest of the paper, we will refer Equation 2 as Hoffman scaling law.
 Coefficients $e$ $a$ $b$ $\alpha$ $\beta$
@@
 TABLE II: Fitted parameter values for Equation 2
-II-B Sparse Scaling Laws
-Sparse models use only a subset of the parameters available in the models.
-As a result, Equation 2 needs to be adjusted and extended.
+II-B Scaling Laws for Sparse Models
+Sparse LLMs use only a subset of the parameters available in the models.
 Several works had proposed scaling laws for sparse transformer models [16, 1, 26].
@@
 In this work we focus on two types of sparsifications: pruning and mixture-of-experts (MoE).
-II-B1 Pruned Transformer Models
-One way to introduce sparsity in transformer models is pruning, where a number of parameters are removed from the architecture and are not used during the forward pass or backpropagation, hence making no contribution to the training process.
-Pruning can be categorized into three main groups [27].
-In structured pruning, entire higher-level units or even whole layers are removed; this process delivers universal speedup but sometimes at the cost of accuracy.
-In unstructured pruning, individual weights are removed from anywhere in the network; usually the least important weights are pruned, with the importance based on certain criteria.
-In semi-structured pruning, one tries to strike a balance between unstructured and structured pruning approaches by enforcing fixed pruning patterns in small regions of the network.
+Figure 2: Comparison between Hoffman scaling law and Abnar scaling law at $0\%$ sparsity.
+II-B1 Pruned Models
+One way to introduce sparsity in transformer models is pruning, where a subset of parameters are removed or masked and are not used during the forward pass or backpropagation, hence making no contribution to the training process.
 Frantar et al.
 [16] studied the scaling behavior of models pruned by unstructured pruning.
-Sparsity $S$ is defined as $S=\frac{\text{Total Parameters - Nonzero Parameters}}{\text{Total Parameters}}$.
-The experimental results with a total of $48$ models trained showed that sparsity introduced by pruning affects only the parameter-dependent term of Equation 2 and has no impact on the data-dependent term.
-The loss was fitted as a function of nonzero parameters $N$, data $D$, and sparsity $S$ as shown in the following equation.
+In their work, the authors defined sparsity, $S=\frac{\text{Total Parameters}-\text{Nonzero Parameters}}{\text{Total Parameters}}$.
+Based on the pretraining results of $48$ models with different sparsity levels, numbers of nonzero parameters and training tokens, the authors claimed that sparsity introduced by pruning affects only the parameter-dependent term of Hoffman scaling law and has minor to no impact on the data-dependent term.
+The authors also claim that increasing sparsity leads to a better pretraining loss.
+Their experimental results also showed that as the number of nonzero parameters and training tokens—in other words, total compute increased the pretraining loss keeps getting better, which holds the claim of Hoffman scaling law [20].
+Using the Hoffman scaling law as the foundation, the authors fitted loss as a function of nonzero parameters $N$, data $D$, and sparsity $S$ as shown in Equation 3.
 $L(N,D,S)=\left(a_{S}(1-S)^{b_{S}}+c_{S}\right).\left(\frac{1}{N}\right)^{b_{N}}+\left(\frac{a_{D}}{D}\right)^{b_{D}}+c$ (3)
-Following the similar approach of [3] to find the optimal values for the coefficients of Equation 3, the best-fitted values for these coefficients are determined empirically through some experiments and fitting the data shown in Table III.
+Following the similar approach of [20] the authors presented the best-fitted values for the scaling law coefficients of Equation 3.
+Table III lists the scaling law parameter values mentioned in [16].
+For the rest of the paper, we will refer to Equation 3 as the Frantar scaling law.
 Coefficients $a_{S}$ $b_{S}$ $c_{S}$ $b_{N}$ $a_{D}$ $b_{D}$ $c$
@@
 TABLE III: Fitted parameter values for Equation 3
-To draw the similarities between Equations 2 and 3, we used the similar symbols in Equation 2 to reformat the coefficients in Equation 3.
+To draw the similarities between Hoffman scaling law in Equation 2 and Frantar scaling law in Equation 3, we used the similar symbols used in Hoffman scaling law to reformat the scaling law coefficients in Frantar scaling law.
 We used $e$ instead of $c$ to capture the lower bound on the loss that represents the inherent stochasticity of the modeling problem.
 We replaced $b_{N}$ and $b_{D}$ with $\alpha$ and $\beta$, respectively, and ${a_{D}}^{b_{D}}$ with $b$.
-The result is the following equation:
+$c_{S}$ is the constant sparsity factor.
+The reformatted version of Frantar scaling law is Equation 4.
 $L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (4)
-Equation 4 helps visualize the transition from the dense scaling law Equation 2 to the sparse scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity.
-Based on the changes in the coefficient representation, the best-fitted values need to be updated correspondingly.
-The updated values for the coefficients in Equation 4 are calculated and listed in Table IV.
+Equation 4 helps visualize the transition from the dense Hoffman scaling law Equation 2 to the sparse Frantar scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity.
+Based on the changes in representation, we recalculate the best-fitted values for the scaling law coefficients in the reformatted Frantar scaling law, and the values are listed in Table IV.
 Coefficients $a_{S}$ $b_{S}$ $c_{S}$ $b$ $\alpha$ $\beta$ $e$
@@
 TABLE IV: Fitted parameter values for Equation 4
-While Equation 3 functionally captures the effect of sparsity on loss, the maximum sparsity used in the experiments was $87.5\%$ [4].
-We have generated an estimated IsoFLOP dataset for the compute budget $1e20$ from the experiments conducted in [5], which included sparsities of more than $87.5\%$.
-When we used the new dataset to plot the scaling behavior of Equation 3, we noticed a spike for smaller nonzero parameters, as shown in Figure 1, where each line connects models with different sparsity and model sizes trained under the same amount of compute budget $1e20$ with nonzero parameters as x-axis and training loss as y-axis.
-Increasing the nonzero parameters results in decreasing the data size.
-However, based on the work in [4, 5], sparsity should provide a smooth performance gain as we increase sparsity.
-This indicates that Equation 3 does not capture the scaling behavior of highly sparse models very well.
-Figure 1: IsoFLOP plot of Equation 3 for different sparsities
-On the other hand, the largest model used in [4] had only $85M$ nonzero parameters.
-Compared with the current work, these models are relatively much smaller.
-This work used the scaling law Equation 2 [3] as the base for the approach.
-For $0\%$ sparsity, or for dense models, the loss should be equal to Equation 2.
-But for the same data set with no sparsity, Equation 3 provides different coefficient values from those of Equation 2, as shown in Figure 2, there is a difference of up to 2.5 in loss.
-Thus, we can conclude that Equation 2 is not a special case of Equation 3 if the sparsity is 0.
-Figure 2: Comparison between Equations 2 and 3 for dense models
+While Equation 4 functionally captures the effect of sparsity on loss, the maximum sparsity used in the experiments was $87.5\%$ [16] while other works [1] show that better performance gain can be achieved by introducing even more sparsity (even at 98% sparsity).
+The largest model used in Frantar scaling law experiments had only $85M$ nonzero parameters which are relatively much smaller.
+Although this work used the Hoffman scaling law as baseline, at $0\%$ sparsity or for dense models with the same dataset, Equation 4 provides different loss predictions than those in Hoffman scaling law with an average MSE (Mean Squared Error) of $3.04$, as shown in Figure 1.
+Thus, we can conclude that Hoffman scaling law is not a special case of Frantar scaling law where sparsity is $0\%$
+Figure 3: Loss prediction of Hoffman scaling law and the proposed scaling law.
 II-B2 MoE Models
-Another way to introduce sparsity is to use the mixture-of-experts method [28] where the feed-forward network for a transformer model is replicated multiple times and each of these copies is called an “expert.” Only a fraction of experts are activated per token, and this number is a fixed value.
+Another way to introduce sparsity is to use the mixture-of-experts (MoE) method [33] where the feed-forward network of a transformer model is replicated multiple times and each of these copies is called an “expert.” Only a fraction of experts are activated per token.
 The tokens are routed to different experts during the pretraining process.
-The size of the experts is usually equal to the size of the feed-forward network.
+(a) Training tokens: $16B$
+(b) Training tokens: $32B$
+(c) Training tokens: $65B$
+Figure 4: As sparsity and the number of nonzero/active parameters increase, the pretraining loss decreases.
+In each figure, the number of non-zero parameters in the models used ranged from 1.3M85M1.3M-85M with varying sparsity levels, and the number of training tokens is mentioned in the subcaption.
 Abnar et al.
 [1] investigated the sparse MoEs through a large-scale empirical study.
-The sparsity $S$ was defined as $S=\frac{E-K}{E}$, where $E$ is the total number of experts and $K$ is the number of active experts.
-For five different compute budgets and seven different sparsities, different models were trained.
-Based on the experimental results, the authors proposed the following scaling law:
+The authors defined sparsity, $S=\frac{E-K}{E}$, where $E$ is the total number of experts and $K$ is the number of active experts.
+For five ($3e19,6e19,1e20,3e20,\text{ and }1e21$) different compute budgets and seven ($0\%,25\%,50\%,75\%,90\%,95\%,\text{ and }98\%$) different sparsities, different models were trained.
+The conclusion they made was that as we increase sparsity along with model size and number of training tokens, the pretraining loss decreases.
+For IsoFLOP experiments, they noticed that each for a specific sparsity level, each model size has a sweet spot.
+Until that spot, the pretraining loss keeps decreasing, and beyond that spot, pretraining loss keeps increasing.
+Based on the experimental results, the authors proposed Equation 5 as their scaling law.
+For the rest of the paper, we will refer to Equation 5 as Abnar scaling law.
 $L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}$ (5)
@@
 TABLE V: Fitted parameter values for Equation 5
-Equation 5 introduced two different terms in the scaling law along with the previous three terms from [3] to capture the impact of sparsity.
+Abnar scaling law introduced two different terms in the scaling law along with the previous three terms from Equation 2 [20] to capture the impact of sparsity.
 The newly introduced terms are interdependent.
-The fifth term in Equation 5 can be interpreted as an active parameter term because the denominator term is a multiplication of $N$ and 1S1-S.
-Because the total number of parameters already includes the active parameters, however, having two different terms for total parameters and active parameters in Equation 5 causes some redundancy in the loss function.
-This work also used Equation 2 as the base.
-But for $0\%$ sparsity we apply the models from [5] to Equations 5 and 2 to generate Figure 3 and notice a difference in loss values.
-This indicates that Equation 2 is not a special case of Equation 5 if the sparsity is 0.
-Figure 3: Comparison between Equations 2 and 5 for dense models
+The fifth term $\frac{d}{(1-S)^{\delta}N^{\gamma}}$ in Equation 5 can be interpreted as an active parameter term because the denominator term is a multiplication of $N$ and 1S1-S.
+As the total number of parameters already includes the active parameters, having two different terms for total parameters and active parameters in Equation 5 can be deemed as redundant.
+While this may help the Abnar scaling law to fit the experimental results better, we later show that this can be represented in a generalized way without introducing redundancy.
+Abnar scaling law also used Hoffman scaling law as the base.
+We compared both of these laws at $0\%$ sparsity using the same model size and data size used in Abnar scaling law experiments.
+The results were off from each other with and average MSE of $1.06$ in Figure 2.
 III Generalizing Scaling Laws
-We observed that the previous work and their scaling laws are based on the number of nonzero or active parameters.
-Dense models use all the parameters, and hence in [3] $N$ was used as the total number of parameters in its scaling law.
-In [4] the nonzero parameters $N$ were considered in its scaling law 3.
-In [5] the authors modified the compute budget $C=6ND$ from [2] to $C=6N_{a}D$, where $N_{a}$ is the number of active parameters.
-These scaling laws do not cover the dense scaling law when the sparsity is 0.
-These observations and limitations motivate us to generalize the scaling laws of [3, 4, 5] into a single representation that can effectively capture the experimental behavior and features observed in these works.
-We want to make sure that the new scaling law works for both dense and sparse models.
-As we increase model size, data size, and compute, we want to get better performance [2, 3, 4, 5].
-Similarly, as we increase the sparsity, we want to obtain performance improvement [4, 5].
-Furthermore, for $0\%$ sparsity or dense models, we need to make sure that our scaling law behaves the same as the original scaling law in Equation 2.
-For simplicity, we use the term active parameters $N$ for active/nonzero parameters in the MoE and pruned models.
+We observed that regardless of the dense or sparse architecture, all the previous works mentioned so far use the number of nonzero or active parameters to measure the total compute.
+Hoffman et al.
+[20] used $C=6ND$ where $N$ is the total number of parameters and for dense models this is the number of nonzero/active parameters.
+Abnar et al.
+[1] used $C=6N_{a}D$ where $N_{a}$ is the number of active parameters.
+Frantar et al.
+[16] used $C=6ND\cdot c_{mul}(S)$ where the authors used dense Hoffman’s compute formula and multiplied it by a sparsity-dependent factor $c_{mul}(S)$.
+We have also observed in the preceding section that the mentioned sparse scaling laws had their limitations and do not replicate the dense scaling law behavior when the sparsity is $0\%$.
+In Section I we have also discussed that due to the architecture-specific nature of existing scaling laws where some scaling laws take total parameters into account for their scaling law and some use the number of nonzero/active parameters, cross-architecture comparison and budget planning become challenging.
+On top of that, different scaling laws use different scaling law coefficients that need to be tuned carefully to fit the early experimental results for better prediction.
+We think that the scaling behavior, regardless of the architecture (dense or sparse), is dependent on the number of active/non-zero parameters and propose a generalized scaling law that can address the earlier-mentioned issues.
+This proposed scaling law should 1) be able to describe the scaling behavior for both dense and sparse models; 2) achieve similar predictions and capture the scaling behavior observed in the experiments of [23, 20, 16, 1] as increasing model size, data size, and compute; and 3) capture the behavior for sparse models as observed in [16, 1] for increasing the sparsity.
+Furthermore, for $0\%$ sparsity or dense models, the scaling law should behave exactly as the Hoffman scaling law.
+Figure 5: Scaling behavior and prediction comparison between Abnar scaling law and our proposed scaling law.
+For simplicity, we use the term “active parameters” ($N$) for active/nonzero parameters in the MoE and pruned models.
 For the dense models, the total number of active parameters is the total number of parameters.
-We use Equation 2 as the base to extend it to a generalized scaling law.
+We use Hoffman scaling law (Equation 2) as the base to propose a generalized scaling law.
 We define the sparsity as follows:
 $S=\frac{\text{Total Parameters - Active Parameters}}{\text{Total Parameters}}$ (6)
-Where $0\leq S<1$ because the active parameters in a model cannot be zero.
-When $S=0$, it means that the total parameters is the same as the active parameters.
-Therefore, $\frac{N}{1-S}$ is the total number of parameters.
-In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$.
-Similar to the approach in [4] we introduce an upper-bound constant sparsity factor $c$, which is caused by the sparsity; but this factor, which contributes to the loss, depends on the sparsity $S$.
+Where $0\leq S<1$ (The number of active parameters in a model cannot be zero and hence, $S\not=1$).
+$\frac{N}{1-S}$ is the total number of parameters.
+In the parameter-dependent second term of Hoffman scaling law in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$.
+Similar to the Frantar et al.
+approach in [16] we introduce an upper-bound constant sparsity factor $c$; but unlike Frantar scaling law, this factor depends on the sparsity $S$.
 So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$.
-For sparsity $S=0$ or dense models, this term is the same as the original parameter-dependent term in Equation 2.
-We have noticed that different approaches adjusted the entropy $e$ in different ways, as shown in Tables II, IV, and V.
-The value of $e$ is impacted by the sparsity.
-We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$, where $\gamma$ is a constant.
-Accumulating these changes, we propose the following scaling law:
+While $e$ corresponds to irreducible loss, empirical fits over $N$ and $D$ treat the constant as an effective offset.
+For sparse models, this offset can vary systematically with sparsity.
+We model this dependence with minimal multiplicative factor $(1-S)^{\gamma}$, preserving the dense limit and power law behavior of scaling laws.
+Accumulating these changes, we propose Equation 7 as our proposed scaling law.
 $L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (7)
 This equation quantifies the impact of sparsity on loss.
-Notice that we did not change the values of the data-dependent term in Equation 2, because the work in [4, 5] has shown experimentally that sparsity has no effect on this term.
-The fitted coefficient values in our proposed Equation 7 are listed in Table VI.
-Since we have used the empirical scaling law in [3] as the base, we simply apply the coefficient values in Table II to Equation 7.
-For the sparsity factor $c$, we use the dataset for the largest model with the largest sparsity in [4] to fit Equation 7 to obtain the value for the sparsity factor $c$.
-The value of $\gamma$ was empirically obtained.
-Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
-Values $1.69$ $406.4$ $410.7$ $93.45$ $0.34$ $0.28$ $10^{-2}$
-TABLE VI: Fitted parameter values for Equation 7
-Notice that Equation 7 is similar to Equation 4; however, the coefficients and the parameter-dependent term are different.
-If the sparsity is 0, Equation 4 is not the same as Equation 2, as shown in Figure 2.
-Lemma: If the sparsity $S$ is 0, Equation 7 is the same as Equation 2.
+Notice that we did not change the values of the data-dependent term ($\frac{b}{D^{\beta}}$) in Equation 2 (Hoffman scaling law), because the work in [16, 1] has shown experimentally that sparsity has no effect on this term.
+Figure 6: IsoFLOP plot of Abnar scaling law (left) and our proposed scaling law (right) for a compute budget of $1e20$ FLOPs.
+Lemma: If sparsity $S$ is 0, Equation 7 is the Hoffman scaling law in Equation 2.
 If the sparsity $S$ is 0, the total number of parameters is equal to the total number of active parameters. From Equation 7 we have
@@
 L(N,D,0)L(N,D)=0L(N,D,0)-L(N,D)=0.
+Figure 7: IsoFLOP plot of Abnar scaling law (left) and our proposed scaling law (right) for a compute budget $3.29e24$ FLOPs of DeepSeek-V3 training.
 IV Performance Results and Discussion
 IV-A Performance Results
-In this section we compare the performance between our proposed scaling law and the previous scaling laws for dense and sparse models.
-We have collected the experimental datasets (number of parameters, number of tokens, and sparsity) used in [3, 4, 5].
+In this section we compare the performance of our proposed scaling law and the previous architecture specific scaling laws for dense and sparse models.
+We have collected the experimental datasets (number of parameters, number of tokens, and sparsity) used in [20, 16, 1] shown in Table VI.
 Hoffman et al.
 [20] provided a list of nine models with number of parameters and training tokens used in their experiments.
-Franta et al.
+Frantar et al.
 [16] provided four different numbers of nonzero parameters, three different numbers of training tokens, and four different sparsity levels with a total of $48$ experimental runs to collect their experimental results.
@@
 [1] used five different numbers of compute budgets, seven different sparsity levels, and a range of model sizes.
-Using the formula $C=6ND$, we calculated their number of training tokens and listed the ranges in Table VII.
-We use these experimental datasets to generate plots for their respective scaling laws, and we use the same datasets to generate plots using our proposed scaling law in order to do a head-to-head comparison.
+Then we use these experimental datasets to generate plots for their respective scaling laws and use the same datasets to generate plots using our proposed scaling law in Equation 7 in order to compare them.
 Scaling Law Parameters Tokens Sparsities(%) Test Models
-Hoffman et al. [3] 400M10T400M{-}10T 8B216.2B8B{-}216.2B - 99
-Frantar et al. [4] 1.3M85M1.3M{-}85M 16B65B16B{-}65B $0,50,75,87.5$ $48$
-Abnar et al. [5] 329M21.2B329M{-}21.2B 15B128B15B{-}128B $0,25,50,75,90,95,98$ $35$
-TABLE VII: Parameters, tokens, sparsity, and number of test models collected from the corresponding works.
-When we compare our proposed scaling law in Equation 7 with the dense scaling law [3] in Equation 2, as we discussed in the lemma in the preceding section, if the sparsity is 0, Equation 7 is the same as Equation 2.
+Hoffman [20] 400M10T400M{-}10T 8B216.2B8B{-}216.2B - 99
+Frantar [16] 1.3M85M1.3M{-}85M 16B65B16B{-}65B $0,50,75,87.5$ $48$
+Abnar [1] 329M21.2B329M{-}21.2B 15B128B15B{-}128B $0,25,50,75,90,95,98$ $35$
+TABLE VI: Parameters, tokens, sparsity, and number of test models collected from the corresponding works.
+Along with drawing the plots, we measure the prediction Mean Square Error (MSE) between our proposed scaling law and the respective scaling law for better visualization of prediction of our proposed scaling law.
+(a) Tuning Loss for Equation 2
+(b) Tuning Loss for Equation 3
+(c) Tuning Loss for Equation 5
+Figure 8: Loss calculated by the scaling law coefficients generated by ytopt and grid search.
+The red lines are the loss using the best scaling law coefficients provided in the respective work, and the blue dots are the loss tuned for different evaluations over time using ytopt.
+First, to compare our proposed scaling law in Equation 7 with Hoffman scaling law in Equation 2 [20], we refer back to the Lemma in section III where we have shown that if the sparsity $S=0$, Equation 7 converts into Equation 2.
+Table VII shows the fitted coefficient values in Equation 7 for the dense models from [20].
 Figure 3 shows the loss vs compute budget using the dataset in [20].
-On the $x$-axis is compute.
-which depends on parameters $N$ and tokens $D$ and hence captures the impact of both ($C=6ND$).
-On the $y$-axis is the calculated loss.
-We see an identical plot, indicating that our proposed scaling law behaves the same as the original scaling law for dense models.
-Figure 4: Comparison between Hoffman’s scaling law in Equation 2 and the proposed scaling law in Equation 7
-We compare the proposed scaling law in Equation 7 with the scaling law [4] in Equation 3.
-We use the 48 models to conduct the experiments in [4] for performance comparison.
-Figures 5 (a) and 5 (b) show that our proposed scaling law replicates the behavior of the scaling law in [4].
-We observe that with the increase of the sparsity the loss decreases.
-However, there is a difference between the loss values for both scaling laws which is similar to Figure 2.
-This is because the models used for the experiments in [4] are relatively much smaller (the maximum number of parameters just 85M) than other works, and the adjusted coefficient values in the proposed scaling law are for larger models, as shown in Figure 4 where the loss decreases with increasing the model size and compute budget.
-Hence we see a relatively larger loss for the much smaller datasets.
-(a) Scaling behavior for Equation 3 with varying compute budget
-(b) Scaling behavior for Equation 7 with varying compute budget
-Figure 5: Comparison between (a) the scaling law in Equation 3 and (b) the proposed scaling law in Equation 7
-We also compare our proposed scaling law in Equation 7 with the scaling law [5] in Equation 5.
-For performance comparison we use the interpolated values of parameters, tokens, and the exact compute budget of the 35 models the authors used to conduct the experiments in [5].
-We have drawn compute budget vs loss plots in Figures 6(a) and 6(b) to compare Equations 5 and 7.
-We observe that with the increase of the sparsity and compute, the loss decreases; and both figures show a similar pattern.
-Increasing the compute budget means an increase in the active parameter size ($N_{a}$) and number of tokens ($D$).
-This indicates that the proposed scaling law shows scaling behavior similar to that in [5].
-(a) Scaling behavior for Equation 5 with varying compute budget
-(b) Scaling behavior for Equation 7 with varying compute budget
-Figure 6: Comparison between (a) the scaling law in Equation 5 and (b) the proposed scaling law in Equation 7
-IV-B Hyperparameter Optimization for Better Coefficients
-Most of the previous scaling laws discussed in this paper used grid search to find the best values for the coefficients.
+On the $x$-axis is compute ($C=6ND$).
+which depends on model size $N$ and tokens $D$ and hence captures the impact of both.
+On the $y$-axis is the calculated loss using both scaling laws.
+We see an identical plot, indicating that our proposed scaling law behaves the same as the original scaling law for dense models capturing the dense scaling behavior where increasing compute results in decreasing pretraining loss.
+Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
+Values $1.69$ $406.4$ $410.7$ 0 $0.34$ $0.28$ $0.01$
+TABLE VII: Fitted scaling law coefficient values of our proposed scaling law for dense models.
+To compare our proposed scaling law with Frantar scaling law, we use the data (sparsity, training steps, number of non-zero parameters and training tokens) collected from the 48 models that were used to conduct the experiments in [16].
+Figure 4 shows that our proposed scaling law captures the almost similar scaling behavior as Frantar scaling law for sparse models as sparsity and number of of nonzero parameters keep increasing.
+The loss prediction is very similar with average $MSE=0.00078$.
+Table VIII shows the fitted parameter values for our proposed scaling law for models from Frantar scaling law experiments [16].
+Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
+Values $0.15$ $86.03$ $7.90$ $29.26$ $0.28$ $0.08$ $2.00$
+TABLE VIII: Fitted parameter values of our proposed scaling law for Frantar-style sparse models.
+We also compared our proposed scaling law with the Abnar scaling law in Equation 5 [1].
+For this comparison we use the interpolated values of parameters, tokens, and the exact compute budget of the 35 models used to conduct the experiments in [1].
+We have used interpolated values to generate the full dataset because even though the authors have shared a list of compute budget and sparsity levels, no such list was shared for model size and training tokens.
+So we used the plots shared in [1] and generated the entire dataset.
+In Figure 5 we have drawn compute budget (FLOPs) vs.
+loss plots to compare the behavior of both scaling laws.
+We observe that with the increase of the sparsity and compute, the loss decreases, and both scaling laws show a similar pattern.
+The average prediction MSE between Abnar scaling law and our proposed scaling law is $0.0005$ which is not significant.
+We present our fitted scaling law coefficient values in Table IX.
+Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
+Values $0.57$ $8.26$ $6324.82$ $3.57$ $0.08$ $0.40$ $1.19$
+TABLE IX: Fitted parameter values of our proposed scaling law for MoE-based models.
+Abnar et al.
+[1] also conducted IsoFLOP experiments with the given compute budget of $1e20$ FLOPs for the MoE-based models.
+In Figure 6 we have shown the IsoFLOP loss prediction comparison between Abnar scaling law and our proposed scaling law, where each line connects models with different sparsity and model sizes trained under the same amount of compute budget $1e20$ with active parameters as x-axis and training loss as y-axis.
+In [1] the authors intentionally cut the plot for active parameters at $4B$ parameters even though the original experiments had active parameters around $26B$.
+For comparison purpose, we mark that $4B$ parameter spot and generate the entire plot in Figure 6.
+For IsoFLOP experiments [1], the authors showed that for fixed sparsity, as we increse number of active parameters, the pretraining loss keeps decreasing until it reaches a sweet spot.
+Beyond that sweet spot, loss starts increasing again.
+As seen from Figure 6, our proposed scaling law can effectively capture this behavior.
+To demonstrate the effectiveness of our scaling law for larger models, we use DeepSeek-V3 [11] which is a MoE-based model with 671B parameters and 37B of which are activated for each token as an example.
+The model was trained with $14.8T$ tokens.
+Using the compute formula $C=6ND$ where $N$ is the number of active parameters and $D$ is the number of training tokens, we measure the total compute, $3.29e24$ used to train DeepSeek-V3.
+We generate an interpolated IsoFLOP dataset using this compute budget where we restricted the number of total parameters to $671B$ parameters.
+Figure 7 shows that our proposed scaling law exhibits a very similar pattern to Abnar scaling law for very large models as well.
+IV-B Hyperparameter Optimization for Better Scaling Law Coefficients
+Most of the previous scaling laws discussed in this paper used grid search to find the best values for the scaling law coefficients.
 While grid search is a commonly used approach, it is still inefficient.
-As an alternative, we have used ytopt [36, 37, 38] for hyperparameter optimization.
+As an alternative, we have used ytopt [39, 41, 42] for scaling law coefficient optimization.
 ytopt is a Bayesian optimization–based autotuner that builds a surrogate model to explore the promising regions to find the optimal values with fewer evaluations.
@@
 Unlike traditional autotuners, ytopt leverages libEnsemble’s [21] asynchronous manager-worker framework to run multiple evaluations in parallel, reducing the one-by-one autotuning time.
-To use ytopt for the optimization process for finding the best coefficients for the previous scaling laws, we use the best coefficients for the scaling laws in Tables II, IV, and V as the baseline and the default values to define the proper search space for each coefficient.
+To use ytopt for the optimization process for finding the best scaling law coefficients for the previous scaling laws, we use the best scaling law coefficients for the scaling laws in Tables II, IV, and V as the baseline and the default values to define the proper search space for each coefficient.
 We carefully define the range for each coefficient for Equations 2, 3, and 5 and then use ytopt to run the optimization process.
-The generated hyperparameter optimization results in Figure 7(a), (b), and (c) illustrate that ytopt identifies the best coefficient configuration (combination of coefficients) over time (blue) with the baseline (red) for the default values and that using ytopt results in lower loss values than the coefficient-based loss values (red, baseline) using grid search.
-(a) Tuning Loss for Equation 2
-(b) Tuning Loss for Equation 3
-(c) Tuning Loss for Equation 5
-Figure 7: Loss calculated by the coefficients generated by ytopt and grid search.
-The red lines are the loss using the best coefficients provided in the respective work, and the blue dots are the loss tuned for different evaluations over time using ytopt.
+The generated hyperparameter optimization results in Figure 8(a), (b), and (c) illustrate that ytopt achieves better loss predictions than grid search.
+Since ytopt is a Bayesian optimization-based approach, it is much faster than a grid search-based approach and is overall a better alternative.
 IV-C Discussion
-In summary, we observe that our proposed scaling law Equation 7 effectively captures the scaling behavior of Equations 2, 3, and 5 for dense and sparse LLMs.
-In Figure 8 the proposed scaling law shows no spike for higher sparsity with the fixed compute budget 1e20 as we observed for Equation 3 in Figure 1.
-This indicates that the proposed scaling law captures the scaling behaviors for higher sparsity.
-For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with the given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt [36, 37].
-In Figure 5 we observe that the range of loss values is different for Equations 3 and 7.
-This difference is mainly because the experiments done for Equation 3 use much smaller model sizes with at most 85M.
-For the scaling law in Equation 5 we can see that the results are almost identical.
-These results further confirm that our proposed scaling law captures the scaling behaviors of the MoE models.
-Figure 8: IsoFLOP plot of the proposed scaling law in Equation 7 for different sparsities
+In summary, we observe that our proposed generalized scaling law Equation 7 effectively captures the scaling behavior of Equations 2, 3, and 5 for dense and sparse LLMs.
+The effectiveness is further justified by the small MSE between the existing scaling law and out proposed scaling law predictions.
+In Figure 6 and 7 we have seen that our proposed scaling law can capture the IsoFLOP behavior of MoE models as well.
+So we can conclude that for a given compute budget, our proposed scaling law can be used to estimate the best model hyperparameters (Model size, Tokens and Compute) for a given sparsity or to identify the optimal sparsity for the given model hyperparameters.
 In this work we have generalized the scaling behaviors of dense and sparse models.
 Our generalized scaling law extends and covers several existing scaling laws.
-Its limitations are related to the sparsity factor $c$ and the coefficient $\gamma$, which need to be further optimized for larger models.
-There are other ways [41, 42] to introduce sparsification in a model.
-Different training approaches [30, 32] may yield different scaling behavior as well.
-Further work can be done to incorporate these scaling laws into our generalized scaling law.
+Its limitations are related to the sparsity factor $c$ and the coefficient $\gamma$, which need to be further optimized for larger LLM models.
 V Conclusions
-Current LLMs result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks.
-However, these models are costly to train.
-As LLMs get larger and more expensive, it is crucial for researchers and developers to estimate the expected performance beforehand.
+As LLMs get larger and more expensive in training, it is crucial for researchers and developers to estimate the expected performance beforehand.
 Scaling laws thus are critical.
-But while architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical.
-Our generalized scaling law covers densely activated, pruned, and MoE models.
+While architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical.
+Our generalized scaling law covers densely activated models, pruned models, and MoE models.
 Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures.
-For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with a given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt.
-Because LLMs inference can be memory-bound and compute-intensive, future work will extend the generalized scaling law in training to not only incorporate other scaling laws [30, 31, 32] but also take into account inference [43], especially recent agentic AI workloads for answering complex questions through reasoning using chain-of-thought prompting.
+Because LLMs inference can be memory-bound and compute-intensive, future work will extend the generalized scaling law in training to not only incorporate other scaling laws [24, 4, 5, 38, 35] but also take into account inference [32], especially recent agentic AI workloads for answering complex questions through reasoning using chain-of-thought prompting.
 Since the behavior of scaling laws varies based on training and inference approaches, these approaches can be incorporated in the generalized scaling law as well.
 Unifying different scaling laws into one single representation lays the foundation of architecture-agnostic representation of the scaling behavior of LLMs.
-As LLM models continue to diversify, this generalized approach will allow us to anticipate the return based on their allocated compute resources.
 Acknowledgments
-This work was supported DOE ASCR RAPIDS2 and OASIS. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.
+This work was supported by DOE ASCR SciDAC RAPIDS and OASIS.
+We acknowledge the Argonne Leadership Computing Facility for use of Polaris under the project EE-ECP.
+This material is based upon work supported by the U.S.
+Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.